# Reinforcement learning-based statistical search strategy for an axion model from flavor

### v1

@@ 4 Comparison with traditional methods
In this section, we show the distribution of terminal states of the quark sector, which is found by the RL. We made 20 agents for each energy scale, and in total, it took 6 days to train the neural networks on a single CPU. The loss functions tend to be minimized, and we also check for an increase in the sum of rewards. Thus, overfittings did not occur.
@@
Our statistical approach with RL is practically more efficient than brute force computation. Let $p$ be the fraction of charge pairs that satisfy $V_{1}=1.0$ in the parameter space generated by the charge combinations. In our learning, we generated the following number of random sets of charges as initial states of each episode:
@@
In those, no terminal state is found. Therefore, it is reasonable to estimate $p<1/\left(8\times 10^{6}\right)$. Let $t$ be the time required to find one terminal state that satisfies $V_{1}=1.75$ using a method other than RL. Then, the total time $T$, which is required to find realistic models with a similar number of charges as those found in this work, can be estimated as $T=2{,}185\cdot t/p$.
We tested various optimization methods of “SciPy”, which is a library of Python, to find $\eta$ satisfying Eq. (5.5). In Table 5, we show the results of actual measurements using the same computer as in the case of RL. It turns out that the time required to find the optimal value of $\eta$ relating to a particular set of FN charges is at least about 0.274 ms. Thus, the total time $T$ is estimated as follows:
@@
We show the distribution for all terminal states in Fig. 1. In addition, the distribution of the top 10 models with the highest intrinsic value is also shown to compare with an aligned number of terminal states at each energy scale. In these, terminal states have high intrinsic values around $M=10^{14},10^{15}\,\mbox{GeV}$.
@@
The distribution of the domain wall number $N_{\mathrm{DW}}$ at each energy scale is shown in Fig. 2. For all scales, the models are located around $N_{\mathrm{DW}}=30$. In addition, Table 6 shows the average of intrinsic values ${\cal V}$ for each $N_{\mathrm{DW}}$ category at each scale. The table shows that the distribution of ${\cal V}$ varies from scale to scale. For example, the intrinsic value tends to be highest in the region of $N_{\mathrm{DW}}\geq 36$ at $M=10^{14},10^{17}\,\mbox{GeV}$, so some cases with small $N_{\mathrm{DW}}$ are not expected to reproduce the renormalized masses. The scale of $M=10^{15}\,\mbox{GeV}$ also tends to have large ${\cal V}$ at about $N_{\mathrm{DW}}=30$ even considering the $1\sigma$ error range. Since the model with the highest intrinsic value among the 156 terminal states is located at $\left(M/\mbox{GeV},N_{\mathrm{DW}}\right)=\left(10^{15},36\right)$, it is still expected to be difficult to reproduce renormalized masses with small $N_{\mathrm{DW}}$. On the other hand, for $M=10^{16}\,\mbox{GeV}$, the category with smaller $N_{\mathrm{DW}}$ tends to have higher ${\cal V}$. Although the analysis with $M=10^{16}\,\mbox{GeV}$ includes the uncertainty due to the relatively large error, it would be possible that there are even smaller $N_{\mathrm{DW}}$ terminal states at $M=10^{16}\,\mbox{GeV}$.
@@
In a previous study Ema:2016ops , an example with $N_{\mathrm{DW}}=26$ was treated as a benchmark for roughly reproducing the flavor structure. By statistical analysis with RL as described above, the lower limit of $N_{\mathrm{DW}}$ reproducing the flavor structure may be around $N_{\mathrm{DW}}=20$. In any case, RL is useful as a flexible method to explore discrete parameters such as FN charges and to consider the constraints from cosmology and so on.
@@
$k=14$ $-2.78\pm 0.13$ $-2.71\pm 0.19$ $\mathbf{-2.61}\pm 0.20$
$k=15$ $-2.89\pm 0.20$ $\mathbf{-2.55}\pm 0.15$ $-2.91\pm 0.24$
$k=16$ $\mathbf{-2.47}\pm 0.27$ $-2.77\pm 0.17$ $-3.12\pm 0.26$
$k=17$ $-2.69\pm 0.16$ $-2.86\pm 0.17$ $\mathbf{-2.44}\pm 0.10$
@@ 5 Exploring the flaxion physics
[…]
In this section, we discuss the flaxion phenomenology. To calculate the flaxion-photon coupling, we should determine the FN charges for both quarks and leptons. Thus, we run some training for the lepton sector. Two benchmark points of the quark sector are shown in Table 7 and Table 8. Benchmark 1 has the highest intrinsic value in the case of $M=10^{15}\,\mbox{GeV}$. On the other hand, benchmark 2 has the highest intrinsic value in the case of $M=10^{16}\,\mbox{GeV}$.
@@
By using those benchmarks, we also make the agent learn the lepton sector. We adopt $N_{\rm ep}=6\times 10^{4}$ and the other hyper-parameters are the same as the learning of the quark sector. We considered renormalized masses of the charged leptons shown in Table 9 and a constraint for the sum of neutrino masses $\Sigma m_{\nu}<85\,\mbox{meV}$ based on the $\Lambda\mathrm{CDM}$ model in Ref. Brieden:2022lsd . Moreover, we also check constraints for the mixing angles based on NuFIT v5.3.
@@
We make 6 agents for each of the benchmarks and need 15 hours on a single CPU for the whole learning process. In the training, we fix the normal ordering of neutrino masses. The number of terminal states found by the agent is summarized in Table 10. For $M=10^{15}\,\mbox{GeV}$, we find 4,732 terminal states, and among them, 23 models satisfy the $3\sigma$ constraints for the mixing angles $\theta_{12},\theta_{13},\theta_{23}$. On the other hand, for $M=10^{16}\,\mbox{GeV}$, we find few terminal states with $V_{2}=0.2$, and no state satisfies all of the constraints for the mixing angles. Thus, similar to the evaluation of renormalized masses, we repeat the learning process with $V_{2}=0.3$ and then narrow down terminal states to $V_{2}=0.2$. We then obtain 6,219 terminal states, and among them, 7 models satisfy the $3\sigma$ constraints for the mixing angles. Finally, 30 models were extracted by considering the mixing constraints, which were chosen as benchmarks for the lepton sector. The details of these benchmarks for the lepton sector are summarized in Appendix A.
@@
The flaxion-photon coupling is given by using the charges obtained from the results of RL. Note that the charges of both quarks and leptons are needed to calculate the coupling. Therefore, we chose the charge pairs for benchmark 1 in Tables 7 and 11 or benchmark 2 in Tables 8 and 12. Fig. 4 shows the effective coupling constant based on the charges of the benchmark models. This mass range will be explored in the future experiment, ex., DMRadio-$\mathrm{m}^{3}$ in Ref. DMRadio:2022pkf . In this figure, only the results at $M=10^{15}\,\mbox{GeV}$ are within the sensitivity region of the DMRadio-$\mathrm{m}^{3}$. However, the calculated values at $M=10^{16}\,\mbox{GeV}$ are also close to the sensitivity region. Therefore, it is expected that the RL with the appropriate value of $V_{2}$ will lead to the discovery of models that overlap in the same sensitivity region in a short time.
@@
Figure 4: Right circles (Red) and left triangles (green) are the results calculated by using the FN charges of the lepton sector corresponding to Benchmarks 1 and 2, respectively. The gray area indicates the region excluded by existing experiments. On the other hand, the blue area indicates the sensitivity of DMRadio-$\mathrm{m}^{3}$ proposed in Ref. DMRadio:2022pkf .
@@ 6 Conclusion
Flavor symmetries give many insights into new physics beyond the Standard Model. Recently, there have been many works analyzing theoretical models with machine learning. Under these circumstances, RL can be a new statistical analysis method to capture features of the models that are difficult to observe in sequential calculations. In particular, RL is useful for solving issues which are difficult within conventional descent approaches.
In this paper, we constructed a new algorithm to search for the model parameters in the minimal axion model from flavor following the RL-based search strategy, where the flavor structure had been analyzed in Ref. Harvey:2021oue and Ref. Nishimura:2023nre . In Sec. 2, we introduced the $U(1)$ flavor model with the FN mechanism, and a pseudo-scalar field arising from the imaginary component of the flavon. Then, in Sec. 3, we designed RL to find $U(1)$ charge assignments of the quark sector and considered renormalizing effects for masses. In Sec. 4, we studied the statistical differences at the four energy scales, i.e., $M=10^{14},10^{15},10^{16},10^{17}$ GeV, by training 20 agents at each scale. Through 6 days of neural network training, the agents found a total of 156 terminal states that had errors within single-digit from the renormalized masses. In addition, we estimated the time required to discover the same number of terminal states without using RL. For various optimization methods, we require the total time of at least 55 days. This indicates that the RL-based search method for discrete parameters is faster than conventional computational methods. We also discussed the phenomenology of the flaxion in Sec. 5, taking the constraints from the dark matter, the isocurvature perturbation and the inflation into account. It turned out that $U(1)$ charge assignments with high intrinsic values are realized around $M=10^{14},10^{15}$ GeV, although the isocurvature constraint puts a lower bound on $M$, i.e., $M\gtrsim 10^{15}$ GeV. Note that the reproducibility of renormalized masses depends on the energy scales. Furthermore, we ran additional training for the lepton sector for the benchmarks of quarks. As a result, we found some terminal states of the leptons that satisfy the constraints of the sum of neutrino masses and mixing angles. From those states, we show the distribution of the flaxion mass and flaxion-photon coupling at $M=10^{15},10^{16}$ GeV, which will be explored in future experiments such as DMRadio-$\mathrm{m}^{3}$ DMRadio:2022pkf . Our proposed RL-based efficient analysis may provide a statistical perspective not only for our discussion of the cosmological axion model but also for other particle phenomenology as well as cosmology.
@@
  • In this work, since only one flavon field is introduced, we are not able to derive the CP-symmetry breaking in the quark sector. One way to improve this situation is to introduce two flavons and consider the phase difference between them. In this case, one of the two flaxions will be regarded as the QCD axion, and the mechanism of mass generation for the other flaxion in the orthogonal direction should be discussed. Alternatively, another way to derive CP-symmetry breaking is to consider the complex Yukawa couplings. In this case, it is nontrivial whether RL can find many terminal states immediately. That is because the parameter space of Yukawa couplings generated in the first step of learning is expanded. Then, it is also useful to discuss how the degree of the FN charge space has been explored by the agent when no terminal state is found in a single training session for fixed Yukawa couplings.
  • We focused on the specific axion model with $U(1)$ flavor symmetry, but our neural network architecture can be applied to other axion models as well as flavors by replacing the input data with the appropriate ones. Furthermore, it would be interesting to verify the axion from other experiments and cosmological observations by using the RL-based strategy, but we leave comprehensive studies for future work.
  • Our architecture of the neural network follows a previous study Nishimura:2023nre . This network has about 10,000 parameters, so in total, one agent has 20,000 parameters to be optimized since two types of neural networks are used in DQN (Q network and target network). In this study, the parameters of the two networks were initialized in the first episode. Since each agent works through a different learning process independently, the parameters of the networks differ for each agent. On the other hand, neural networks are known to exhibit the phenomenon of “generalization”. This is the ability to make appropriate classifications or predictions for unknown data that have never been observed. Our networks output optimal actions even for sets of FN charges that have not been inputted before. This attempt itself is established based on the generalization ability. However, in the sense that the generalization is only valid for certain fixed Yukawa couplings, this study does not fully make good use of the unique strengths of neural networks. Thus, preparing a larger-scale architecture is a stepping stone to realizing a more generalized search strategy by RL. If we construct neural networks that can output the optimal action for arbitrary Yukawa couplings, the learning itself can be done only once, and further reduction of computational time is expected.

### v2

@@ 4 Comparison with traditional methods
In this section, we show the distribution of terminal states for the quark sector, which is found by the RL. We made 20 agents for each energy scale, and in total, it took six days to train the neural networks on a single CPU. The loss functions tend to be minimized, and we also check for an increase in the sum of rewards. Thus, overfittings did not occur.
@@
The predicted flavor structure is strongly dependent on the assignments of $U(1)$ charges. In terms of intrinsic values, the evaluation of each model fluctuates dramatically with changes in discrete parameters. Therefore, the usual gradient descent method cannot be used to obtain good charges that reproduce the flavor structure, and we must rely on brute force calculations or random searches. With this background, our statistical approach with RL is practically more efficient than brute force computation. Let $p$ be the fraction of charge pairs that satisfy $V_{1}=1.0$ in the parameter space generated by the charge combinations. In our learning, we generated the following number of random sets of charges as initial states of each episode:
@@
Since no terminal state is found in those sets, it is reasonable to estimate $p<1/\left(8\times 10^{6}\right)$. Let $t$ be the time required to find one terminal state that satisfies $V_{1}=1.75$ using a method other than RL. Then, the total time $T$, which is required to find realistic models with a similar number of charges as those found in this work, can be estimated as $T=2{,}185\cdot t/p$.
We tested various optimization methods of “SciPy”, which is a library of Python, to find $\eta$ satisfying Eq. (3.12). In Table 5, we show the results of actual measurements using the same computer as in the case of RL. It turns out that 0.274 ms is required at least to find the optimal value of $\eta$ such that maximize the intrinsic value $\cal V$} under a particular set of FN charges. Thus, the total time $T$ is estimated as follows:
@@
We show the distribution for all terminal states in Fig. 1. In addition, the distribution of the top 10 models with the highest intrinsic value is also shown to compare with an aligned number of terminal states at each energy scale. In Fig. 1, the quartile ranges are represented by boxes, and the orange lines represent the medians. The upper and lower ends of the error bars refer to the maximum and minimum intrinsic values of the models, respectively. Taking into account all terminal states, there is no significant difference in the median and interquartile range of intrinsic values for each energy scale. This suggests that the training of the agents is not biased at each scale and stable learning is realized. In fact, the behavior of the loss function and rewards also showed no occurrence of overfitting.
On the other hand, the number of terminal states found by RL varies across different energy scales. Once sufficiently trained, agents tend to seek out terminal states as quickly as possible within their environment in order to avoid negative rewards (punishments) associated with reductions in intrinsic value. However, if no terminal state exists near the randomly assigned initial charge, the agents fail to reach a solution and simply proceed to the next episode. Therefore, while the learning process remains stable, the variation in the number of terminal states results in differing ratios of solutions corresponding to these terminal states. Moreover, terminal states have high intrinsic values in the case of $M=10^{14}$ and $10^{15}\,\mbox{GeV}$. Taken together, it is inferred that there are relatively many solutions that can reproduce the experimentally confirmed flavor structure of quarks at $M=10^{14},10^{15}\,\mbox{GeV}$. The two perspectives presented above enable us to consider the differences between the energy scales. However, future large-scale searches will be necessary to provide additional statistical support for each piece of evidence.
@@
The distribution of the domain wall number $N_{\mathrm{DW}}$ at each energy scale is shown in Fig. 2. For all scales, the models are located around $N_{\mathrm{DW}}=30$. In addition, Table 6 shows the average of intrinsic values $\cal V$} for each $N_{\mathrm{DW}}$ category at each scale. The table shows that the distribution of $\cal V$} varies from scale to scale. For example, the intrinsic value tends to be highest in the region of $N_{\mathrm{DW}}\geq 36$ at $M=10^{14},10^{17}\,\mbox{GeV}$, so some cases with small $N_{\mathrm{DW}}$ are not expected to reproduce the renormalized masses. The scale of $M=10^{15}\,\mbox{GeV}$ also tends to have large $\cal V$} at about $N_{\mathrm{DW}}=30$ even considering the $1\sigma$ error range. Since the model with the highest intrinsic value among the 156 terminal states is located at $\left(M/\mbox{GeV},N_{\mathrm{DW}}\right)=\left(10^{15},36\right)$, it is still expected to be difficult to reproduce renormalized masses with small $N_{\mathrm{DW}}$. On the other hand, for $M=10^{16}\,\mbox{GeV}$, the category with smaller $N_{\mathrm{DW}}$ tends to have higher $\cal V$}. Although the analysis with $M=10^{16}\,\mbox{GeV}$ includes the uncertainty due to the relatively large error, it would be possible that there are even smaller $N_{\mathrm{DW}}$ terminal states at $M=10^{16}\,\mbox{GeV}$. The potential for discovering nontrivial relationships through RL is promising, but more precise investigation is required in the future.
@@
In a previous study Ema:2016ops , an example with $N_{\mathrm{DW}}=26$ was treated as a benchmark for roughly reproducing the flavor structure. By statistical analysis with RL as described above, the lower limit of $N_{\mathrm{DW}}$ reproducing the flavor structure may be around $N_{\mathrm{DW}}=20$, summarized in Fig. 2 and Table 6. In any case, RL is useful as a flexible method to explore discrete parameters such as FN charges and to consider the constraints from cosmology and so on.
@@
$k=14$ $2.78\pm 0.13$ $2.71\pm 0.19$ $\mathbf{-2.61}\pm 0.20$
$k=15$ $2.89\pm 0.20$ $\mathbf{-2.55}\pm 0.15$ $2.91\pm 0.24$
$k=16$ $\mathbf{-2.47}\pm 0.27$ $2.77\pm 0.17$ $3.12\pm 0.26$
$k=17$ $2.69\pm 0.16$ $2.86\pm 0.17$ $\mathbf{-2.44}\pm 0.10$
@@ 5 Exploring the flaxion physics
Each of the models found by the agents takes a different charge configuration. Depending on their charges, the flaxion decay constant $f_{a}$ as well as the misalignment angle $\theta_{i}$ given by Eq. (2.29) are determined by results discovered by the agent. Then, based on the results of RL, we can set the upper bound for the inflationary scale $H_{\mathrm{inf}}$ through Eq. (2.30). Fig. 3 shows the correlation between inflationary scales and the misalignment angles for each model. The colors of the data classify energy scales. The blue, red, green, and orange mean $M=10^{14},\ 10^{15},\ 10^{16},$ and $10^{17}\,\mbox{GeV}$ respectively. The gray shaded region is excluded by Eq. (2.32). Therefore, we can state that $M\gtrsim 10^{15}\,\mbox{GeV}$ models are favored by cosmology.
In this section, we discuss the flaxion phenomenology. To calculate the flaxion-photon coupling, we should determine the FN charges for both quarks and leptons. Thus, we run some training for the lepton sector. Two benchmark points of the quark sector are shown in Table 7 and Table 8. Benchmark 1 has the highest intrinsic value in the case of $M=10^{15}\,\mbox{GeV}$. On the other hand, benchmark 2 has the highest intrinsic value in the case of $M=10^{16}\,\mbox{GeV}$. By definition, models with such high intrinsic values tend to successfully reproduce the renormalized masses and mixings. Therefore, these benchmarks can be regarded as providing the most suitable Lagrangian for exploring flaxion physics. In the quark sector, since the Yukawa couplings are treated as ordinary continuous parameters, it is possible to reproduce the flavor structure with even greater accuracy by optimizing the parameters using conventional methods, such as Markov Chain Monte Carlo. However, this approach requires a long computational time due to the large number of coupling constants involved.
@@
We make the agent learn the lepton sector for these benchmarks. We adopt $N_{\rm ep}=6\times 10^{4}$ and the other hyper-parameters are the same as the learning of the quark sector. We considered renormalized masses of the charged leptons shown in Table 9 and a constraint for the sum of neutrino masses $\Sigma m_{\nu}<85\,\mbox{meV}$ based on the $\Lambda\mathrm{CDM}$ model in Ref. Brieden:2022lsd . Moreover, we also check constraints for the mixing angles based on NuFIT v5.3.
@@
We make six agents for each of the benchmarks and need 15 hours on a single CPU for the whole learning process. In the training, we choose the normal ordering of neutrino masses. The number of terminal states found by the agent is summarized in Table 10. For $M=10^{15}\,\mbox{GeV}$, we find 4,732 terminal states, and among them, 23 models satisfy the $3\sigma$ constraints for the mixing angles $\theta_{12},\theta_{13},\theta_{23}$. On the other hand, for $M=10^{16}\,\mbox{GeV}$, we find few terminal states with $V_{2}=0.2$, and no state satisfies all of the constraints for the mixing angles. Thus, similar to the evaluation of renormalized masses, we repeat the learning process with $V_{2}=0.3$ and then narrow down terminal states to $V_{2}=0.2$. We then obtain 6,219 terminal states, and among them, 7 models satisfy the $3\sigma$ constraints for the mixing angles. Finally, 30 models were extracted by considering the mixing constraints, which were chosen as benchmarks for the lepton sector. The details of these benchmarks for the lepton sector are summarized in Appendix A.
@@
The flaxion-photon coupling is given by using the charges obtained from the results of RL. Note that the charges of both quarks and leptons are needed to calculate the coupling. Therefore, we adopt two benchmark charge assignments: benchmark 1 with the quark and lepton charges listed in Tables 7 and 11, respectively, and benchmark 2 with the corresponding charges given in Tables 8 and 12. Fig. 4 shows the effective coupling constant based on the charges of the benchmark models. This mass range will be explored in the future experiment, ex., DMRadio-$\mathrm{m}^{3}$ in Ref. DMRadio:2022pkf . In this figure, only the results at $M=10^{15}\,\mbox{GeV}$ are within the sensitivity region of the DMRadio-$\mathrm{m}^{3}$. However, the calculated values at $M=10^{16}\,\mbox{GeV}$ are also close to the sensitivity region. Therefore, it is expected that the RL with the appropriate value of $V_{2}$ will lead to the discovery of models that overlap in the same sensitivity region in a short time.
@@
Figure 4: Right circles (Red) and left triangles (green) are the results calculated by using the FN charges of the lepton sector corresponding to Benchmarks 1 and 2, respectively. The gray area indicates the region excluded by existing experiments PhysRevLett.104.041301 ; ADMX:2018ogs ; ADMX:2019uok ; Salemi:2021gck ; PhysRevLett.127.261803 ; Noordhuis:2022ljw ; Benabou:2025jcv summarized in Ref. AxionLimits . On the other hand, the blue area indicates the sensitivity of DMRadio-$\mathrm{m}^{3}$ proposed in Ref. DMRadio:2022pkf .
@@ 6 Conclusion
Flavor symmetries give many insights into new physics beyond the Standard Model. Recently, numerous studies have analyzed theoretical models using machine learning techniques. Under these circumstances, RL can be a new statistical analysis method to capture features of the models that are difficult to observe in sequential calculations. In particular, RL is useful for solving issues which are difficult within conventional descent approaches.
In this paper, we constructed a new algorithm to search for the model parameters in the minimal axion model from flavor, utilizing the RL-based search strategy. The flavor structure has been previously analyzed in Refs. Nishimura:2023nre and Harvey:2021oue . In Sec. 2, we introduced the $U(1)$ flavor model with the FN mechanism, and a pseudo-scalar field arising from the imaginary component of the flavon. Then, in Sec. 3, we developed an RL to find $U(1)$ charge assignments of the quark sector and considered renormalizing effects on masses. In Sec. 4, we studied the statistical differences at the four energy scales, i.e., $M=10^{14},10^{15},10^{16},10^{17}$ GeV, by training 20 agents at each scale. Through six days of neural network training, the agents found a total of 156 terminal states that had errors within single-digit from the renormalized masses. In addition, we estimated the time required to discover the same number of terminal states without using RL. Various conventional optimization methods would require at least 55 days to perform the same task. This indicates that the RL-based search method for discrete parameters is faster than conventional computational methods. We also discussed the phenomenology of the flaxion in Sec. 5, taking the constraints from the dark matter, the isocurvature perturbation and the inflation into account. It turned out that $U(1)$ charge assignments with high intrinsic values are realized around $M=10^{14},10^{15}$ GeV, although the isocurvature constraint puts a lower bound on $M$, i.e., $M\gtrsim 10^{15}$ GeV. Note that the reproducibility of renormalized masses depends on the energy scales. Furthermore, we ran additional training for the lepton sector for the benchmarks of quarks. As a result, we found some terminal states for the leptons that satisfy the constraints on the sum of neutrino masses and mixing angles. From those states, we showed the distribution of the flaxion mass and flaxion-photon coupling at $M=10^{15},10^{16}$ GeV, which will be explored in future experiments such as $\mathrm{m}^{3} DMRadio$:2022pkf . Our proposed RL-based efficient analysis may provide a statistical perspective not only for our discussion of the cosmological axion model but also for other areas of particle phenomenology and cosmology.
@@
  • In this work, since only one flavon field is introduced, we are not able to derive the CP-symmetry breaking in the quark sector. One way to improve this situation is to introduce two flavons and consider the phase difference between them. In this case, one of the two flaxions will be regarded as the QCD axion, and the mechanism for mass generation of the other flaxion in the orthogonal direction should be discussed. Alternatively, another way to derive CP-symmetry breaking is to consider the complex Yukawa couplings. In this case, it is nontrivial whether the RL can find many terminal states immediately. That is because the parameter space of Yukawa couplings generated in the first step of learning is expanded. Then, it is also useful to discuss how the degree of the FN charge space has been explored by the agent when no terminal state is found during a single training session with fixed Yukawa couplings.
  • We focused on the specific axion model with $U(1)$ flavor symmetry, but our neural network architecture can be applied to other axion models as well as flavors by replacing the input data with the appropriate ones. Furthermore, it would be interesting to verify the axion through other experiments and cosmological observations using the RL-based strategy, but we leave comprehensive studies for future works.
  • Our architecture of the neural network follows a previous study Nishimura:2023nre . This network has about 10,000 parameters, so in total, one agent has 20,000 parameters to be optimized since two types of neural networks are used in DQN (Q network and target network). In this study, the parameters of both networks were initialized in the first episode. Since each agent works through a different learning process independently, the parameters of the networks differ for each agent. On the other hand, neural networks are known to exhibit the phenomenon of “generalization”. This is the ability to make appropriate classifications or predictions for unknown data that have never been observed. Our networks output optimal actions even for sets of FN charges that have not been inputted before. This attempt itself is established based on the generalization ability. However, since the generalization holds only for specific fixed values of the Yukawa couplings, this study does not fully exploit the distinctive capabilities of neural networks. Thus, preparing a larger-scale architecture is a stepping stone to realizing a more generalized search strategy by RL. If we construct neural networks that can output the optimal action for arbitrary Yukawa couplings, the learning itself can be done only once, and further reduction of computational time is expected.